# Ophthalmic Biomarker Detection Using Ensembled Vision Transformers and Knowledge Distillation

### v2

@@ 4 CONCLUSION
In this work, we outlined the methodology of our Ophthalmic Biomarker Detection study. We also presented the underlying motivation for pipeline design decisions. We find that Vision Transformer (ViT) models have begun to consistently outperform their Convolutional Neural Network (CNN) counterparts. Furthermore, we find that k-fold cross-validation and model ensembling continue to be effective means to utilize the entire dataset and to improve the generalization of predictions.
@@ 5 ACKNOWLEDGEMENT
We would like to extend our sincere gratitude to Dr. S.M. Rezwan Hussain, a distinguished ophthalmologist at the Eye Department, Combined Military Hospital(CMH), Dhaka. Bangladesh, for his invaluable insights and expertise regarding biomarker classification according to their spatial extent.

### v3

@@ 6 CONCLUSION
In this work, we outlined the methodology for our study on Ophthalmic Biomarker Detection and presented the underlying motivations for pipeline design decisions. Our findings indicate that Vision Transformer (ViT) models have begun to consistently outperform their Convolutional Neural Network (CNN) counterparts. Additionally, we observed that k-fold cross-validation and model ensembling are effective techniques for leveraging the entire dataset and improving generalization. Finally, utilizing the abundance of unlabeled data through knowledge distillation proves to be an efficient approach for enhancing model performance. For future work, we plan to explore our pipeline’s generalizability to patient data collected from diverse sources, and interpretability analysis to improve trustworthiness.
@@ 7 ACKNOWLEDGEMENT
We would like to extend our sincere gratitude to Dr. S.M. Rezwan Hussain, a distinguished ophthalmologist at the Eye Department, Combined Military Hospital (CMH), Dhaka, Bangladesh, for his invaluable insights and expertise regarding biomarker classification according to their spatial extent.